Consider the matrices


A = 3 0 2 , B = -5 1 1 , C = 1 1 1

4 -6 3 0 3 0 2 3 -1

-2 1 8 7 6 2 3 -5 -7


Verify the following expressions (where possible and give reasons)

(i) A + (B + C) = (A + B) + C and A(BC) = (AB)C.

(ii) (a − b)C = aC + bC and a(B − C) = aB − aC, where a = −2, b = 3 .

(iii) (AT)T = A and (A − B)T = A T− BT



1
Expert's answer
2021-05-17T18:40:02-0400

(i)


"A+(B+C)=\begin{pmatrix}
 3 & 0 & 2 \\
 4 & -6 & 3 \\
 -2 & 1 & 8
\end{pmatrix}"

"+\bigg(\begin{pmatrix}
 -5 & 1 & 1 \\
 0 & 3 & 0 \\
 7 & 6 & 2
\end{pmatrix}+\begin{pmatrix}
 1 & 1 & 1\\
 2 & 3 & -1 \\
 3 & -5 & -7
\end{pmatrix}\bigg)"




"=\begin{pmatrix}
 3 & 0 & 2 \\
 4 & -6 & 3 \\
 -2 & 1 & 8
\end{pmatrix}+\begin{pmatrix}
 -4 & 2 & 2 \\
 2 & 6 & -1 \\
 10 & 1 & -5
\end{pmatrix}"

"=\begin{pmatrix}
 -1 & 2 & 4 \\
 6 & 0 & 2 \\
 8 & 2 & 3
\end{pmatrix}"

"(A+B)+C=\bigg(\begin{pmatrix}
 3 & 0 & 2 \\
 4 & -6 & 3 \\
 -2 & 1 & 8
\end{pmatrix}"


"+\begin{pmatrix}
 -5 & 1 & 1 \\
 0 & 3 & 0 \\
 7 & 6 & 2
\end{pmatrix}\bigg)+\begin{pmatrix}
 1 & 1 & 1\\
 2 & 3 & -1 \\
 3 & -5 & -7
\end{pmatrix}"

"=\begin{pmatrix}
 -2 & 1 & 3 \\
 4 & -3 & 3 \\
 5 & 7 & 10
\end{pmatrix}+\begin{pmatrix}
 1 & 1 & 1\\
 2 & 3 & -1 \\
 3 & -5 & -7
\end{pmatrix}"

"=\begin{pmatrix}
 -1 & 2 & 4 \\
 6 & 0 & 2 \\
 8 & 2 & 3
\end{pmatrix}"

"A+(B+C)=\begin{pmatrix}
 -1 & 2 & 4 \\
 6 & 0 & 2 \\
 8 & 2 & 3
\end{pmatrix}=(A+B)+C"


"BC=\begin{pmatrix}
 -5 & 1 & 1 \\
 0 & 3 & 0 \\
 7 & 6 & 2
\end{pmatrix}\begin{pmatrix}
 1 & 1 & 1\\
 2 & 3 & -1 \\
 3 & -5 & -7
\end{pmatrix}"


"=\begin{pmatrix}
 0 & -7 & -13 \\
 6 & 9 & -3 \\
 25 & 15 & -13
\end{pmatrix}"

"A\big(BC\big)=\begin{pmatrix}
 3 & 0 & 2 \\
 4 & -6 & 3 \\
 -2 & 1 & 8
\end{pmatrix}\begin{pmatrix}
 0 & -7 & -13 \\
 6 & 9 & -3 \\
 25 & 15 & -13
\end{pmatrix}"

"=\begin{pmatrix}
 50 & 9 & -65 \\
 39 & -37 & -73 \\
 206 & 143 &-81
\end{pmatrix}"

"AB=\begin{pmatrix}
 3 & 0 & 2 \\
 4 & -6 & 3 \\
 -2 & 1 & 8
\end{pmatrix}\begin{pmatrix}
 -5 & 1 & 1 \\
 0 & 3 & 0 \\
 7 & 6& 2
\end{pmatrix}"

"=\begin{pmatrix}
 -1 & 15 & 7 \\
 1 & 4 &10 \\
 66 & 49 & 14
\end{pmatrix}"

"\big(AB\big)C=\begin{pmatrix}
 -1 & 15& 7 \\
 1 & 4 & 10 \\
 66 & 49 & 14
\end{pmatrix}\begin{pmatrix}
 1 & 1 & 1\\
 2 & 3 & -1 \\
 3 & -5 & -7
\end{pmatrix}"

"=\begin{pmatrix}
 50 & 9 & -65\\
 39 & -37 & -73 \\
 206 & 143 & -81
\end{pmatrix}"

"A\big(BC\big)=\begin{pmatrix}
 50 & 9 & -65\\
 39 & -37 & -73 \\
 206 & 143 & -81
\end{pmatrix}=A\big(BC\big)"

(ii)


"a=-2, b=3"

"(a-b)C=(-2-3)\begin{pmatrix}
 1 & 1 & 1\\
 2 & 3 & -1 \\
 3 & -5 & -7
\end{pmatrix}"

"=-5\begin{pmatrix}
 1 & 1 & 1\\
 2 & 3 & -1 \\
 3 & -5 & -7
\end{pmatrix}=\begin{pmatrix}
 -5 & -5 & -5\\
 -10 & -15 & 5 \\
 -15 & 25 & 35\end{pmatrix}"

"aC=-2\begin{pmatrix}
 1 & 1 & 1\\
 2 & 3 & -1 \\
 3 & -5 & -7
\end{pmatrix}=\begin{pmatrix}
 -2 & -2 & -2\\
 -4 & -6 & 2 \\
 -6 & 10 & 14
\end{pmatrix}"

"bC=3\begin{pmatrix}
 1 & 1 & 1\\
 2 & 3 & -1 \\
 3 & -5 & -7
\end{pmatrix}=\begin{pmatrix}
 3 & 3 & 3\\
 6 & 9 & -3 \\
 9 & -15 & -21
\end{pmatrix}"

"aC-bC=\begin{pmatrix}
 -2 & -2 & -2\\
 -4 & -6 & 2 \\
 -6 & 10 & 14
\end{pmatrix}-\begin{pmatrix}
 3 & 3 & 3\\
 6 & 9 & -3 \\
 9 & -15 & -21
\end{pmatrix}"

"=\begin{pmatrix}
 -5 & -5 & -5\\
 -10 & -15 & 5 \\
 -15 & 25 & 35\end{pmatrix}"

"(a-b)C=\begin{pmatrix}
 -5 & -5 & -5\\
 -10 & -15 & 5 \\
 -15 & 25 & 35\end{pmatrix}=aC-bC"


"B-C=\begin{pmatrix}
 -5 & 1 & 1 \\
 0 & 3 & 0 \\
 7 & 6 & 2
\end{pmatrix}-\begin{pmatrix}
 1 & 1 & 1\\
 2 & 3 & -1 \\
 3 & -5 & -7
\end{pmatrix}"

"=\begin{pmatrix}
 -6 & 0 & 0\\
 -2 & 0 & 1 \\
 4 & 11& 9
\end{pmatrix}"

"a(B-C)=-2\begin{pmatrix}
 -6 & 0 & 0\\
 -2 & 0 & 1 \\
 4 & 11& 9
\end{pmatrix}=\begin{pmatrix}
 12 & 0 & 0\\
 4 & 0 & -2 \\
 -8 & -22 & -18
\end{pmatrix}"

"aB=-2\begin{pmatrix}
 -5 & 1 & 1 \\
 0 & 3 & 0 \\
 7 & 6 & 2
\end{pmatrix}=\begin{pmatrix}
 10 & -2 & -2 \\
 0 & -6 & 0 \\
 -14 & -12 &-4
\end{pmatrix}"

"aB-aC=\begin{pmatrix}
 10 & -2 & -2 \\
 0 & -6 & 0 \\
 -14 & -12 &-4
\end{pmatrix}"

"-\begin{pmatrix}
 -2 & -2 & -2\\
 -4 & -6 & 2 \\
 -6 & 10 & 14
\end{pmatrix}=\begin{pmatrix}
 12 & 0 & 0\\
 4 & 0 & -2 \\
 -8 & -22& -18
\end{pmatrix}"

"a(B-C)=\begin{pmatrix}
 12 & 0 & 0\\
 4 & 0 & -2 \\
 -8 & -22& -18
\end{pmatrix}=aB-aC"

(iii)


"A^T=\begin{pmatrix}
 3 & 0 & 2 \\
 4 & -6 & 3 \\
 -2 & 1 & 8
\end{pmatrix}^T=\begin{pmatrix}
 3 & 4 & -2 \\
 0 & -6 & 1 \\
 2 & 3 & 8
\end{pmatrix}"

"(A^T)^T=\begin{pmatrix}
 3 & 4 & -2 \\
 0 & -6 & 1 \\
 2 & 3 & 8
\end{pmatrix}^T=\begin{pmatrix}
 3 & 0 & 2 \\
 4 & -6 & 3 \\
 -2 & 1 & 8
\end{pmatrix}=A"

"(A^T)^T=\begin{pmatrix}
 3 & 0 & 2 \\
 4 & -6 & 3 \\
 -2 & 1 & 8
\end{pmatrix}=A"


"A-B=\begin{pmatrix}
 3 & 0 & 2 \\
 4 & -6 & 3 \\
 -2 & 1 & 8
\end{pmatrix}-\begin{pmatrix}
 -5 & 1 & 1 \\
 0 & 3 & 0 \\
 7 & 6 & 2
\end{pmatrix}"

"=\begin{pmatrix}
 8 & -1 & 1 \\
 4 & -9 & 3 \\
 -9 & -5 & 6
\end{pmatrix}"

"(A-B)^T=\begin{pmatrix}
 8 & -1 & 1 \\
 4 & -9 & 3 \\
 -9 & -5 & 6
\end{pmatrix}^T=\begin{pmatrix}
 8 & 4 & -9 \\
 -1 & -9 & -5 \\
 1 & 3 & 6
\end{pmatrix}"

"B^T=\begin{pmatrix}
 -5 & 1 & 1 \\
 0 & 3 & 0 \\
 7 & 6 & 2
\end{pmatrix}^T=\begin{pmatrix}
 -5 & 0 & 7 \\
 1 & 3 & 6 \\
 1 & 0& 2
\end{pmatrix}"

"A^T-B^T=\begin{pmatrix}
 3 & 4 & -2 \\
 0 & -6 & 1 \\
 2 & 3 & 8
\end{pmatrix}-\begin{pmatrix}
 -5 & 0 & 7 \\
 1 & 3 & 6 \\
 1 & 0& 2
\end{pmatrix}"

"=\begin{pmatrix}
 8 & 4 & -9 \\
 -1 & -9 & -5 \\
 1 & 3 &6
\end{pmatrix}"

"(A-B)^T=\begin{pmatrix}
 8 & 4 & -9 \\
 -1 & -9 & -5 \\
 1 & 3 & 6
\end{pmatrix}=A^T-B^T"